# The self-interaction effects on the Kerr black hole superradiance and their observational implications

### v1

@@ IV.2 GW frequency shift
Figure 4: The frequency shift of the bound state calculated with numerical method in the Kerr approximation. When $\alpha$ is small, the numerical result agrees well with the flat-spacetime estimation, Eq. (33).
@@
where $M_{211}$ is the $\,\left|211\right\rangle$ state cloud mass and we consider the attractive self-interaction here (the minus sign). Therefore, we can compare this flat-spacetime estimation with the result calculated using the method described in Sec. II. The numerically computed result in the Kerr approximation is shown in red line in Fig. 4. It can be seen that, when $\alpha$ is significantly smaller than unity, there is a good agreement between the numerical results and the analytical estimations. The red curve approximately follows a trend scaled by $\alpha$ to the fourth power in Fig. 4. However, in the regime of $\alpha\gtrsim 0.2$, the analytical result Eq. (33) might not be a good estimation (dashed line part) and the non-relativistic approximation becomes inapplicable, as has been studied in Ref. [50]. Similarly, in the non-relativistic limit, the self-gravity correction can be estimated by
@@
Moreover, the boson cloud annihilation generates the GWs with frequency twice the boson frequency. During the boson cloud evolution, such as the depletion of cloud mass through annihilation, the bound-state boson frequencies can vary with time, which results in a non-zero GW frequency drift in the signals sourced from the cloud. Those imprints on GWs are perhaps observable in continuous wave searches. The frequency of GWs from the boson cloud when considering the self-interaction is
@@
where ${f}_{\rm GW,\ free}$ is GW frequency for the boson cloud without self-interaction effect. As an example, in the non-relativistic approximation, the GW frequency drift due to self-interaction is determined by
$\dot{f}_{\rm GW}=$ $-3.8\times 10^{-5}\frac{\alpha^{5}}{M_{\text{BH}}^{2}}\left(\frac{1}{f_{a}^{2}}\right)\dot{M}_{211}\ ,$ (36)
where the change of BH mass is neglected. That corresponds to the stage when the bound state has saturated. After the cloud saturation, the cloud slowly loses its mass by radiating continuous GWs. Note that the estimation Eq. (36) for the GW frequency drift can be improved if we use numerical approach to investigate the frequency eigenvalues of boson clouds, especially in the large-$\alpha$ regime.
@@ V Observational consequences on superradiance with self-interaction
In this section, we study further the self-interaction’s observational consequences on the gravitational atom and boson cloud. We focus on two direct consequences of superradiant cloud formation. First, since the BH will spin down during the superradiance process, the spin measurements on the astrophysical BHs are usually used to detect or constrain the bosons with particular mass. Also, the boson cloud will emit GWs through varied processes, which can be used to explore the property of boson cloud and gravitational atom.
@@ V.1 Relaxing constraints on scalar boson models
The direct implication of the self-interaction on the formation of boson cloud is that the cloud growth ceases so long as the occupation number of bound state reaches a critical value $\widetilde{N}_{c}$, which is much smaller than the maximum occupation number ${N}_{c}$ calculated when no self-interaction effect is included. That also means the boson cloud of particular state can extract less angular momenta from the BH than expected. The BH spin-down would pause earlier before its spin became smaller than the respective critical bound for the superradiance of the free field particle. Moreover, the BH spin-down rate is also smaller because of the self-interaction effect. What we can interpret from these is the fast-spinning astrophysical BH can live long with a high dimensionless spin even though the superradiance once happens, if we assume the bosons have a moderate quartic self-interaction. That means our analysis about the self-interaction effect on the superradiance process can relax the existed constraints on the boson mass range obtained through the astrophysical BH spin measurements.
We take the boson with mass $m_{a}=2\times 10^{-13}\leavevmode\nobreak\ \mathrm{eV}$ and an isolated astrophysical BH $M_{\text{BH}}=60\leavevmode\nobreak\ {\rm M_{\odot}}$ as the example. Assuming the initial dimensionless spin $\chi_{i}=0.95$ after the BH’s birth, we concentrate on the $\,\left|211\right\rangle$ state cloud growth, which has the fastest superradiance rate among all states. The BH is isolated from other celestial objects, so the angular momentum can only be transferred from BH to the cloud and keep conservative. When there is not self-interaction in bosons, the $\,\left|211\right\rangle$ cloud will take $\sim 180$ e-folding time to grow to maximum number approximately, and the spin difference is $\sim\mathcal{O}(0.1)$ for the BH spin-down. However, with the existence of self-interaction, the BH spin difference after the cloud stops to grow decreases. We consider the quartic self-interaction with $f_{a}=10^{19},10^{18}\leavevmode\nobreak\ \mathrm{GeV}$ and Planck energy scale, respectively, and calculate the corresponding BH spin difference. The results are shown in the left panel of Table 1. We find that the difference in BH spin after the $\,\left|211\right\rangle$ state cloud growth ceases is smaller than $4\times 10^{-5}$, when the parameter $f_{a}\lesssim 10^{18}\leavevmode\nobreak\ \mathrm{GeV}$. In that regime, the occupation number is suppressed by a factor of order $10^{4}$ at least. That indicates the BH spin-down is very insignificant, different from the situation for the free field boson.
$\chi_{i}=0.95,\,M_{\text{BH}}=60\leavevmode\nobreak\ {\rm M_{\odot}},\,m_{a}=2\times 10^{-13}\leavevmode\nobreak\ \mathrm{eV}$
$f_{a}$ $M_{\mathrm{Pl}}$ $10^{19}\leavevmode\nobreak\ \mathrm{GeV}$ $10^{18}\leavevmode\nobreak\ \mathrm{GeV}$
$\widetilde{N}_{c}$ $\sim 1.5\times 10^{77}$ $\sim 10^{77}$ $\sim 10^{75}$
$\dfrac{\widetilde{N}_{c}}{{N}_{c}}$ $\sim 1\%$ $5.7\times 10^{-3}$ $5.7\times 10^{-5}$
$\varDelta\chi$ $5.1\times 10^{-3}$ $3.4\times 10^{-3}$ $3.4\times 10^{-5}$
@@
Figure 5: The final BH spin after the cloud growth for the primary BH in GW190517. We find when $f_{a}\lesssim 1.91\times 10^{20}\leavevmode\nobreak\ \mathrm{GeV}$, the BH can exist with spin covered by the $95\%$ credible region interpreted from observation data, and when $f_{a}\lesssim 1.6\times 10^{20}\leavevmode\nobreak\ \mathrm{GeV}$, that BH can exist having a large spin contained in the $68\%$ credible interval interpreted from GW data.
@@
Currently, LIGO/Virgo/KAGRA (LVK) Collaborations have reported more than one hundred GW events, most of which are BBH merger. The BH spin measurement through GW detection are used to constrain the boson mass range [41]. In particular, GW190412 [80] and GW190517 give the most stringent constraints on the scalar boson mass among all events in GWTC-2 [41]. If there is long enough time for the bound state boson to grow to a macroscopic cloud before the BHs merge, each BH dimensionless spin should not exceed the critical spin determined by the boson mass and the fastest growing state. That is basically why the BH spin measurement through GW detection can provide the evidence on excluding the boson mass range. However, the individual BHs can retain a large spin until they merge even though the boson of the appropriate mass exists, as long as boson particles have the self-interaction. We take the event GW190517 as the example and assume the initial BH spin when it is formed is equal to the mean value of respective BH spin interpreted from GW data (e.g. for the primary BH in GW190517, it is approximately 0.95). We compute the BH spin difference at the time the $\,\left|211\right\rangle$ cloud growth saturates as the function of the self-interaction scale $f_{a}$. Fig. 5 shows the final BH spin after the cloud growth for the primary BH in the binary, and we can figure out that when $f_{a}\lesssim 1.91\times 10^{20}\leavevmode\nobreak\ \mathrm{GeV}$, the BH can exist with spin covered by the $95\%$ credible region interpreted from GW data. Moreover, when $f_{a}\lesssim 1.6\times 10^{20}\leavevmode\nobreak\ \mathrm{GeV}$, that BH can exist having a large spin contained in the $68\%$ credible interval. Similarly for the primary BH in the event GW190412, these scales are $3.3\times 10^{20}\leavevmode\nobreak\ \mathrm{GeV}$ and $2.6\times 10^{20}\leavevmode\nobreak\ \mathrm{GeV}$, corresponding to $95\%$ and $68\%$ credible intervals, respectively. In total, we can draw a conclusion that if there exists self-interaction with $f_{a}\leq 10^{20}\leavevmode\nobreak\ \mathrm{GeV}$ for the boson mass of $m_{a}=2\times 10^{-13}\leavevmode\nobreak\ \mathrm{eV}$ , the events GW190412 and GW190517 can not be used to constrain the boson model in that mass range via their BH spin measurements. This self-interaction scale is larger than the scale requirement derived in Ref. [41], in which the scale $f_{a}$ shall be greater than $10^{14}\leavevmode\nobreak\ \mathrm{GeV}$ so as to safely neglect the bosenova effect in the boson cloud of both events. That implies the self-interaction effects in the analysis of our work are more common.
@@ V.2 Implication for detecting gravitational atom
The evolution scenario of the BH parameters and the GW signals sourced from the boson cloud through its annihilation could be very different when the bosons have the self-interaction. For a gravitational atom, the mass and spin of BH are interesting observables in the detection, as well as the strength of GWs emitted by its boson cloud. In the prior section, we have studied the consequences of the self-interaction effects. These consequences can be summarized in the following. (1) The BH will have a larger spin and mass after the cloud growth terminates, because the self-interaction brings an extra limitation on the occupation number of bound state and the boson cloud growth will cease when its occupation number reaches $\widetilde{N}_{c}$, which depends on the self-interaction scale $f_{a}$ and BH mass. (2) The self-interaction of boson cloud alters the cloud frequency by an amount dependent on the self-interaction strength and the cloud mass. That is because the attractive self-interaction can bind the macroscopic cloud more tightly, and vice versa. The cloud mass decreases slowly when the cloud annihilates and radiates GW. Therefore the GW frequency change during that evolution, and this drift is possibly observable in the detection. (3) The lighter boson cloud and smaller occupation number reduce the GW emission power from the cloud. The peak GW strain decreases for the GW signals originated from bound state cloud through bosons annihilation, while the GW emission timescale increases to a certain extent accordingly. In order to clearly show the influence of the self-interaction, we numerically compute the evolutions of the BH mass and spin, the occupation number of boson cloud in the gravitational atom, the cloud’s GW strain at the detectors, the GW emission power, and the GW frequency, as the function of time starting from the birth of gravitational atom. We set the benchmark parameters as $M_{\text{BH, init}}=25\leavevmode\nobreak\ {\rm M_{\odot}},\ \chi_{i}=0.95,\ m_{a}=10^{-12}\leavevmode\nobreak\ \mathrm{eV},\ f_{a}=2.5\times 10^{18}\leavevmode\nobreak\ \mathrm{GeV}$.
@@
Figure 6: Evolution examples of BH mass, BH spin, ultralight boson cloud occupation number and its mass, cloud’s radiated GW strain, GW emission power, and GW frequency. The red lines in each subfigure show the evolution of these parameters as the function of time starting from the birth of the gravitational atom. The BH has an initial mass of $M_{\text{BH, init}}=25\leavevmode\nobreak\ {\rm M_{\odot}}$ and an initial spin of $\chi_{i}=0.95$. The self-interaction scale is $f_{a}=2.5\times 10^{18}\leavevmode\nobreak\ \mathrm{GeV}$. To make a comparison, the gray dashed lines in each subfigure show the evolution of the parameters if we assume no self-interaction between bosons in the cloud. We assume the luminosity distance of gravitational atom $D_{L}=10\leavevmode\nobreak\ \text{kpc}$ to compute the GW strain at the detectors. The times in the horizontal axes are all rescaled logarithmically.
@@
We demonstrate the evolutions of the gravitational atom’s parameters in Fig. 6. We assume the luminosity distance of gravitational atom $D_{L}=10\leavevmode\nobreak\ \text{kpc}$ to compute the GW strain at the detectors. Note that the times in the horizontal axes are all rescaled logarithmically. We only consider the boson cloud of fastest growing state for simplicity and assume the attractive self-interaction. The red lines in Fig. 6 show the parameters’ evolution after the gravitational atom is born. As a contrast, we also plot the evolution of those parameters for the free boson in gray dashed lines, in which the self-interaction scale $f_{a}$ is set to be infinite. We can see from the plots the cloud growth ceased before the original saturation and this turn-off happened earlier. The BH mass and spin only decreased slightly and then nearly remained unchanged. The magnitude of boson cloud was much smaller comparing to the boson of free field. As a result, the GW strain was suppressed to a large extent. Although the boson cloud was radiating GWs, its occupation number did not decline after the time $t>3\leavevmode\nobreak\ \text{years}$. The reason for that is the BH with a large spin can continuously replenish the cloud through superradiance when the bosons in the cloud annihilated into graviton. On the whole, the (peak) GW strain became smaller because of the self-interaction. The smaller GW strain and the frequency drift (not shown in the plots) make it more difficult to discover the signals in the GW searches.
@@
VI CONCLUSION
The successful detection of gravitational waves has opened new avenues and provided powerful probes for studying ultralight bosonic fields and dark matter. Traditional dark matter searches often rely on its electromagnetic or other interactions with ordinary matter. However, if dark matter consists of ultralight bosonic fields, it can spontaneously form and grow into macroscopic, dense boson clouds around rotating black holes without requiring any interactions beyond gravity. The gravitational atom formed by a rotating black hole and its surrounding boson cloud exhibits various observable gravitational wave signatures, including modifications to gravitational wave emissions from compact binaries [27, 20, 23] and continuous gravitational wave radiation from the boson cloud itself [28, 29, 30, 31, 32, 33, 34, 35].
On the other hand, if the ultralight bosons constituting the cloud possess self-interactions, these can significantly affect the formation, evolution, and saturation of the boson cloud, the evolution of the black hole, and the gravitational wave signals from the gravitational atom. When the self-interaction is sufficiently strong, both the final state of the black hole and the mass of the boson cloud can differ substantially from the non-interacting case.
In this work, we have studied the formation, growth rate, saturation, and bound state energy levels of self-interacting boson clouds during the black hole superradiance process within the full Kerr background, while also accounting for the corresponding evolution of the black hole’s mass and spin. Previous studies on dark matter detection or constraints via superradiance have largely overlooked the self-interaction effects of boson clouds. Here, we numerically solve for the superradiant growth rate of boson clouds with self-interactions around Kerr black holes and quantitatively evaluate how the strength of scalar boson self-interactions influences the growth rate.

### v2

@@ IV.2 GW frequency shift
Figure 6: The frequency shift of the bound state calculated with numerical method within the Kerr framework. When $\alpha$ is small, the numerical result agrees well with the flat-spacetime estimation, Eq. (34).
@@
where $M_{211}$ is the $\left|211\right\rangle$ state cloud mass and we consider the attractive self-interaction here (the minus sign). Therefore, we can compare this flat-spacetime estimation (34) with the result calculated using the method described in Sec. II. The numerically computed result within the Kerr framework is shown in red line in Fig. 6. It can be seen that, when $\alpha$ is significantly smaller than unity, there is a good agreement between the numerical results and the analytical estimations. The red line (numerical result) in Fig. 6 approximately follows a trend scaled by $\alpha$ to the fourth power. However, in the regime of $\alpha\gtrsim 0.2$, the analytical result Eq. (34) might not be a good estimation (dashed line part) and the non-relativistic approximation used in the analytical method becomes inapplicable, as has been studied in Ref. [51]. Similarly, in the non-relativistic limit [^5] In the large-$\alpha$ regime, Ref. [85] gives a more careful treatment for the self-gravity effects., the self-gravity correction can be estimated by
@@
Moreover, the boson cloud annihilation generates the GWs with frequency twice the boson frequency. During the boson cloud evolution, such as the depletion of cloud mass through annihilation, the boson frequency can vary with time, which results in a non-zero GW frequency drift in the signals sourced from the cloud. Those imprints on GWs are perhaps observable in continuous wave searches. The frequency of GWs from the boson cloud when considering the self-interaction is
@@
where ${f}_{\rm GW,\ free}$ is GW frequency for the boson cloud without self-interaction effect. The additional term in Eq. (36) contributes to the non-zero GW frequency drift, because the time derivative of cloud mass $\dot{M}_{211}$ is non-zero. The depletion rate of cloud is largest when the occupation number reaches the upper bound. That corresponds to the stage when the boson cloud has saturated. After the cloud saturation, the cloud slowly loses its mass by radiating continuous GWs. Last but not the least, from the perspective of GW detection, the frequency shift induced by the self-interaction effects could be degenerate with other parameters, such as boson mass and BH mass. Based on the previous analysis, the amplitude of frequency shift brought by the self-interaction depends on both the self-interaction scale and the occupation number of cloud. Thus, in contrast to the frequency shift caused by other parameters, the frequency shift induced by the self-interaction effects varies differently during the evolution of boson cloud. This characteristic can be conducive to distinguishing the signal of self-interacting boson cloud in the detection.
@@ V Observational consequences on superradiance with self-interaction
In this section, we study further the self-interaction’s observational consequences on the gravitational atom and boson cloud. We focus on two direct consequences of superradiant cloud formation. First, since the BH will spin down during the superradiance process, the spin measurements on the astrophysical BHs are usually used to detect or constrain the bosons with particular mass. Also, the boson cloud will emit GWs through varied processes, which can be used to explore the properties of the boson cloud and the gravitational atom.
@@ V.1 Relaxing constraints on scalar boson models
The direct implication of the self-interaction on the formation of boson cloud is that the cloud growth ceases so long as the occupation number of bound state reaches a critical value $\widetilde{N}_{c}$, which is much smaller than the maximum occupation number ${N}_{c}$ calculated when no self-interaction effect is included. That also means the boson cloud of particular state can extract less angular momenta from the BH than expected. The BH spin-down would pause earlier before its spin reached the corresponding critical bound for the superradiance of the free field particle. Moreover, the BH spin-down rate is also smaller because of the self-interaction effect. What we can interpret from these is the fast-spinning astrophysical BH can live long with a high dimensionless spin even though the superradiance once happens, if we assume the bosons have a moderate quartic self-interaction. That means our analysis about the self-interaction effect on the superradiance process can relax the existed constraints on the boson mass range obtained through the astrophysical BH spin measurements.
$\chi_{i}=0.95,\,M_{\text{BH}}=40\penalty 10000\ {\rm M_{\odot}},\,m_{a}=2\times 10^{-13}\penalty 10000\ \mathrm{eV}$
$f_{a}$ $M_{\mathrm{Pl}}$ $10^{19}\penalty 10000\ \mathrm{GeV}$ $10^{18}\penalty 10000\ \mathrm{GeV}$
$\widetilde{N}_{c}$ $2.6\times 10^{76}$ $\sim 10^{76}$ $\sim 10^{74}$
$\dfrac{\widetilde{N}_{c}}{{N}_{c}}$ 2.82.8\ 1.9 ‰ $1.9\times 10^{-5}$
$\varDelta\chi$ $2.0\times 10^{-3}$ $1.3\times 10^{-3}$ $1.3\times 10^{-5}$
@@
$\chi_{i}=0.95,\,m_{a}=2\times 10^{-13}\penalty 10000\ \mathrm{eV},\,f_{a}=10^{19}\penalty 10000\ \mathrm{GeV}$
$M_{\text{BH}}\penalty 10000\ [{\rm M_{\odot}}]$ 40 60 100
$\widetilde{N}_{c}$ $\sim 10^{76}$ $\sim 10^{77}$ $7.8\times 10^{77}$
$\dfrac{\widetilde{N}_{c}}{{N}_{c}}$ 1.9 ‰ 5.7 ‰ 2.3 %
$\varDelta\chi$ $1.3\times 10^{-3}$ $3.4\times 10^{-3}$ $9.3\times 10^{-3}$
Table 2: Same as Table 1, but for different BH mass $M_{\text{BH}}$ (with a fixed self-interaction strength).
For the boson cloud growth with the self-interaction, the exponential growth rate of the dominant state is determined by not only the fine-structure constant $\alpha$, the boson mass $m_{a}$ and BH spin $\chi$, but also the self-interaction scale $f_{a}$ and the occupation number of the state $N(t)$. These functional dependences have already been explored in Sec. III. We can define the normalized growth rate of boson cloud $\omega_{\text{I}}={\text{Im}(\omega_{a})}/{m_{a}}$ , where the growth rate is normalized by the boson mass and becomes dimensionless in natural units. As the occupation number $N(t)$ increases during the cloud growth, the self-interaction effects become more significant, leading to progressive suppression of the growth rate $\omega_{\text{I}}$ (cf. Fig. 3). Letting the growth rate approaches zero, i.e.
$\omega_{\text{I}}(N,\ f_{a},\ \alpha,\ \chi)=0\,\,$, (37)
we can obtain the critical value of occupation number, $\widetilde{N}_{c}$. When $N=\widetilde{N}_{c}$, the cloud growth ceases, and we can estimate the BH spin difference through the conservation of angular momentum:
$\varDelta\chi=\frac{m\widetilde{N}_{c}}{M_{\text{BH}}^{2}}.$ (38)
Note that $m$ is the magnetic quantum number of bound-state boson.
In contrast, if there is not scalar self-interaction, the boson cloud could saturated and the final BH spin in that stage was generally determined by the superradiance condition (4). Thus, the BH spin difference for free-field case is
$\varDelta\chi_{\rm free}=\left|\chi_{f}-\chi_{i}\right|,$ (39)
where $\chi_{f}$ is BH spin after boson cloud formation and $\chi_{i}$ is initial BH spin.
We take the boson with mass $m_{a}=2\times 10^{-13}\penalty 10000\ \mathrm{eV}$ and an isolated astrophysical BH $M_{\text{BH}}=40\penalty 10000\ {\rm M_{\odot}}$ as the example. Assuming the initial dimensionless spin $\chi_{i}=0.95$ after the BH’s birth, we concentrate on the $\left|211\right\rangle$ state cloud growth, which has the fastest superradiance rate among all states. The BH is isolated from other celestial objects, so the angular momentum can only be transferred from BH to the cloud and keep conservative. When there is not self-interaction in bosons, the $\left|211\right\rangle$ cloud will take $\sim 180$ e-folding time to grow to maximum number approximately, and the spin difference is $\sim\mathcal{O}(0.1)$ for the BH spin-down. However, with the existence of self-interaction, the BH spin difference after the cloud stops to grow will decrease. We consider the quartic self-interaction with $f_{a}=10^{19},10^{18}\penalty 10000\ \mathrm{GeV}$ and Planck energy scale, respectively, and calculate the corresponding BH spin difference. The results are shown in Table 1. We find that the difference in BH spin after the $\left|211\right\rangle$ state cloud growth ceases is smaller than $2\times 10^{-5}$, when the parameter $f_{a}\lesssim 10^{18}\penalty 10000\ \mathrm{GeV}$. In that regime, the occupation number is suppressed by a factor of order $10^{4}$ at least. That indicates the BH spin-down is very insignificant, different from the situation for the free field boson. Similarly, we investigate the BH spin difference for BH systems with initial masses of $40$, $60$, and $100\penalty 10000\ {\rm M_{\odot}}$, respectively, by fixing the self-interaction strength. These results are shown in Table 2. We can see from Table 2 that for a BH with larger initial mass, its spin-down due to the formation of (self-interacting) boson cloud is also larger. Specifically, for the benchmark parameters in Table 2, the spin difference of a $100\penalty 10000\ {\rm M_{\odot}}$ BH is $\varDelta\chi\sim$ 0.01. This value is still 40 times smaller than the result for free-field case.
@@
GW190412 GW190517
primary BH mass $\rm M_{\odot}$}] $27.7$ $39.2$
secondary BH mass $\rm M_{\odot}$}] 9.0 24.0
mass of the remnant BH $\rm M_{\odot}$}] 35.6 60.1
primary BH spin 0.44 0.95
spin of the remnant BH 0.66 0.87
Table 3: The parameters of the BBH systems in GW190412 and GW190517.
@@
Figure 7: The final BH spin after the cloud growth for the primary BH in GW190517. We find when $f_{a}\lesssim 1.91\times 10^{20}\penalty 10000\ \mathrm{GeV}$, the BH can exist with spin covered by the $95\%$ credible region interpreted from observation data, and when $f_{a}\lesssim 1.6\times 10^{20}\penalty 10000\ \mathrm{GeV}$, that BH can exist having a large spin contained in the $68\%$ credible interval interpreted from GW data.
@@
Currently, LIGO/Virgo/KAGRA Collaborations have reported more than one hundred GW events, most of which are BBH merger. The BH spin measurement through GW detection are used to constrain the boson mass range [42]. In particular, GW190412 [86] and GW190517 give the most stringent constraints on the scalar boson mass among all events in GWTC-2 [42]. The parameters of the BHs in these two events are listed in Table 3. If there is long enough time for the bound state boson to grow to a macroscopic cloud before the BHs merge, each BH dimensionless spin should not exceed the critical spin determined by the boson mass and the fastest growing state. That is basically why the BH spin measurement through GW detection can provide the evidence on excluding the boson mass range. However, the individual BHs can retain a large spin until they merge even though the boson of the appropriate mass exists, as long as boson particles have the self-interaction. We take the event GW190517 as the example and assume the initial BH spin when it is formed is equal to the mean value of respective BH spin interpreted from GW data (e.g. for the primary BH in GW190517, it is approximately 0.95). We compute the BH spin difference at the time the $\left|211\right\rangle$ cloud growth saturates as the function of the self-interaction scale $f_{a}$. Fig. 7 shows the final BH spin after the cloud growth for the primary BH in the binary, and we can figure out that when $f_{a}\lesssim 1.91\times 10^{20}\penalty 10000\ \mathrm{GeV}$, the BH can exist with spin covered by the $95\%$ credible region interpreted from GW data. Moreover, when $f_{a}\lesssim 1.6\times 10^{20}\penalty 10000\ \mathrm{GeV}$, that BH can exist having a large spin contained in the $68\%$ credible interval. Similarly for the primary BH in the event GW190412, these scales are $3.3\times 10^{20}\penalty 10000\ \mathrm{GeV}$ and $2.6\times 10^{20}\penalty 10000\ \mathrm{GeV}$, corresponding to $95\%$ and $68\%$ credible intervals, respectively. In total, we can draw a conclusion that if there exists self-interaction with $f_{a}\leq 10^{20}\penalty 10000\ \mathrm{GeV}$ for the boson mass of $m_{a}=2\times 10^{-13}\penalty 10000\ \mathrm{eV}$, the events GW190412 and GW190517 can not be used to constrain the boson model in that mass range via their BH spin measurements. This self-interaction scale is larger than the scale requirement derived in Ref. [42], in which the scale $f_{a}$ shall be greater than $10^{14}\penalty 10000\ \mathrm{GeV}$ so as to safely neglect the bosenova effect in the boson cloud of both events. That implies the self-interaction effects in the analysis of our work are more common.
@@ V.2 Implication for detecting gravitational atom
The evolution scenario of the BH parameters and the GW signals sourced from the boson cloud through its annihilation could be very different when the bosons have the self-interaction. For a gravitational atom, the mass and spin of BH are interesting observables in the detection, as well as the strength of GWs emitted by its boson cloud. In the prior section, we have studied the consequences of the self-interaction effects. These consequences can be summarized in the following. (1) The BH will have a larger spin and mass after the cloud growth terminates, because the self-interaction brings an extra limitation on the occupation number of bound state and the boson cloud growth will cease when its occupation number reaches $\widetilde{N}_{c}$, which depends on the self-interaction scale $f_{a}$ and BH mass. (2) The self-interaction of boson cloud alters the cloud frequency by an amount dependent on the self-interaction strength and the cloud mass. That is because the attractive self-interaction can bind the macroscopic cloud more tightly, and vice versa. The cloud mass decreases slowly when the cloud annihilates and radiates GW. Therefore the GW frequency change during that evolution, and this drift is possibly observable in the detection. (3) The lighter boson cloud and smaller occupation number reduce the GW emission power from the cloud. The peak GW strain decreases for the GW signals originated from bound state cloud through bosons annihilation, while the GW emission timescale increases to a certain extent accordingly. In order to clearly show the influence of the self-interaction, we numerically compute the evolutions of the BH mass and spin, the occupation number of boson cloud in the gravitational atom, the cloud’s GW strain at the detectors, the GW emission power, and the GW frequency, as the function of time starting from the birth of gravitational atom. We set the benchmark parameters as $M_{\text{BH, init}}=25\penalty 10000\ {\rm M_{\odot}},\ \chi_{i}=0.95,\ m_{a}=10^{-12}\penalty 10000\ \mathrm{eV\ and}\ f_{a}=2.5\times 10^{18}\penalty 10000\ \mathrm{GeV}$.
@@
Figure 8: Evolution examples of BH mass, BH spin, ultralight boson cloud occupation number and its mass, cloud’s radiated GW strain, GW emission power, and GW frequency. The red lines in each subfigure show the evolution of these parameters as the function of time starting from the birth of the gravitational atom. The BH has an initial mass of $M_{\text{BH, init}}=25\penalty 10000\ {\rm M_{\odot}}$ and an initial spin of $\chi_{i}=0.95$. The self-interaction scale is $f_{a}=2.5\times 10^{18}\penalty 10000\ \mathrm{GeV}$. To make a comparison, the gray dashed lines in each subfigure show the evolution of the parameters if we assume no self-interaction between bosons in the cloud. We assume the luminosity distance of gravitational atom $D_{L}=10\penalty 10000\ \text{kpc}$ to compute the GW strain at the detectors. The times in the horizontal axes are all rescaled logarithmically.
To clearly show the relation between the GW detection and the gravitational atom system, it is worth to explicate the methods used to derive the GW emission power and GW strain detected. Firstly, in the annihilation process of bosons, the time derivative of boson cloud occupation number is governed by
$\frac{\text{d}N(t)}{\text{d}t}=-\Gamma_{a}N^{2}(t).$ (40)
Here, the minus sign represents the depletion of boson cloud occupation number and this differential equation is equivalent to the Eq. (33) in Sec. IV, the evolution function of occupation number. Then, the emission energy for GWs radiated by boson cloud equals the energy loss of cloud, which means
$P_{\rm GW}=P_{\rm loss}=-\frac{\text{d}N(t)}{\text{d}t}m_{a}.$ (41)
Here we approximate the energy of bound-state boson by $m_{a}$. From the perspective of this work, we focus on how the self-interaction effects change the maximum occupation number and lead to the significant reduction of peak GW strain. We neglect the self-interaction corrections to the annihilation rate and the energy of scalar boson. When calculating the emission power of GWs from boson cloud, those corrections are expected to be subdominant because the amplification of occupation number is largely suppressed. Finally, the (peak) GW strain at the detector can be estimated as
$h\simeq\sqrt{\frac{4P_{\rm GW}}{5\left(2\pi f_{\rm GW}\right)^{2}D_{L}^{2}}},$ (42)
where $D_{L}$ is the luminosity distance of gravitational atom and the prefactor $4/5$ is decided by the configuration of detector.
@@
We demonstrate the evolutions of the gravitational atom’s parameters in Fig. 8. We assume the luminosity distance of gravitational atom $D_{L}=10\penalty 10000\ \text{kpc}$ to compute the GW strain at the detectors. Note that the times in the horizontal axes are all rescaled logarithmically. We only consider the boson cloud of fastest growing state for simplicity and assume the attractive self-interaction. The red lines in Fig. 8 show the parameters’ evolution after the gravitational atom is born. As a contrast, we also plot the evolution of those parameters for the free boson in gray dashed lines, in which the self-interaction scale $f_{a}$ is set to be infinite. We can see from the plots the cloud growth ceased before the original saturation and this turn-off happened earlier. The BH mass and spin only decreased slightly and then nearly remained unchanged. The magnitude of boson cloud was much smaller comparing to the boson of free field. As a result, the GW strain was suppressed to a large extent. Although the boson cloud was radiating GWs, its occupation number did not decline after the time $t>3\penalty 10000\ \text{years}$. The reason for that is the BH with a large spin can continuously replenish the cloud through superradiance when the bosons in the cloud annihilated. On the whole, the (peak) GW strain became smaller because of the self-interaction. The smaller GW strain and the frequency drift (not shown in the plots) make it more difficult to discover the signals in the GW searches.
@@
VI Conclusions and discussions
The successful detection of GWs has opened new avenues and provided powerful probes for studying ultralight bosonic fields and dark matter. Traditional dark matter searches often rely on its electromagnetic or other interactions with ordinary matter. However, if dark matter consists of ultralight bosonic fields, it can spontaneously form and grow into macroscopic, dense boson clouds around rotating BHs without requiring any interactions beyond gravity. The gravitational atom formed by a rotating BH and its surrounding boson cloud exhibits various observable GW signatures, including modifications to GW emissions from compact binaries [28, 21, 24] and continuous GW radiation from the boson cloud itself [29, 30, 31, 32, 33, 34, 35, 36].
Apart from these, whether the ultralight bosons possess self-interactions can significantly affect: (i) the formation, saturation, and evolution of the boson cloud; (ii) the evolution of the BH; and (iii) the GW signals from the gravitational atom. When the self-interaction is sufficiently strong, both the final state of the BH and the mass of the boson cloud can differ substantially from the non self-interacting case.
In this work, we have studied the formation, growth rate, saturation, and bound state energy levels of self-interacting boson clouds during the BH superradiance process within the full Kerr background, while also accounting for the corresponding evolution of the BH’s mass and spin. Previous studies on dark matter detection or constraints via superradiance have largely overlooked the self-interaction effects of boson clouds. Here, we numerically solve for the superradiant growth rate of boson clouds with self-interactions around Kerr BHs and quantitatively evaluate how the strength of scalar boson self-interactions influences the growth rate.